# thanks: corresponding author

### v1

@@ Abstract
In this work, we establish the equivalence between Support Vector Machine (SVM) kernels and quantum Green’s functions. Drawing on the analogy between margin maximization in SVMs and action extremization in Lagrangian mechanics, we show that many standard kernels correspond naturally to Green’s functions and that this correspondence arises from the inversion of physical operators. We further demonstrate how positive semi-definiteness, which is essential for valid SVM kernels, aligns with the spectral properties that ensure well-defined Green’s functions.
We employ the Kernel Polynomial Method (KPM) to create custom kernels for cases where the commonly employed kernels don’t lead to a convergence. These custom kernels approximate the desired Green’s functions. We furthermore demonstrate numerically on examples taken from physical problems, such as electrical conductivity, scattering amplitudes, photonic crystals, and energy levels of anharmonic oscillators, that selecting kernel functions that mirror the mathematical form of the associated Green’s function can significantly enhance the predictive accuracy of machine learning models.
1 Introduction, Motivation and Background
Support Vector Machines (SVMs) have established themselves as a foundational technique in machine learning due to their robustness in classification and regression tasks [1, 2]. Central to their efficacy is the kernel trick, which implicitly maps input data into high-dimensional feature spaces, enabling the separation of complex, nonlinear patterns [3]. Concurrently, quantum mechanics employs Green’s functions as fundamental tools to describe the propagation of particles and the response of systems to external perturbations [4].
1.1 Purpose and Contributions of the Paper
This paper provides a rigorous exploration of the mathematical and physical equivalence between SVM kernel functions and quantum Green’s functions. Through comprehensive studies on various quantum systems, we demonstrate that selecting kernel functions mirroring the mathematical form of the Green’s function associated with a particular physical quantity can significantly improve the accuracy and reliability of machine learning models.
In particular, our contributions include:
  • Establishing a mathematical correspondence between common SVM kernels and quantum Green’s functions.
  • Providing physical interpretations of key theorems, such as Mercer’s theorem and the representer theorem, within the context of quantum mechanics.
  • Introducing the Kernel Polynomial Method (KPM) as a tool for designing custom kernels aligned with specific quantum systems.
  • Demonstrating practical applications, including computing electrical conductivity using SVM regression, scattering amplitude calculations, anharmonic oscillator energy level predictions, and photonic crystal modeling.
  • Analyzing the limitations of certain kernels, particularly the sigmoid kernel, and proposing modifications to ensure their correspondence with Green’s functions.
1.2 Related works
@@
Among the most recent approaches, we have Dean et al. [8] who use kernels to describe quantum correlations of $N$ non-interacting spinless fermions in their ground state. Then, they provide a method to compute the kernel in terms of the Green’s function for the corresponding single particle Schrödinger equation. On the other hand, Gin et al. [9] for example use deep learning to learn the Green’s function kernels with the purpose of solving non-linear differential equations. Li et al. [10] do the same for linear partial differential equations.
The literature in the field is limited, and mostly constrained to applications in mathematics. As far as we can tell, all the previous studies were concerned with identifying equivalence relations between single kernels and Green’s functions. In this paper, for the first time, we provide a collective treatment of the equivalence between Green’s functions and kernel methods with applications in machine learning. We also propose a way to create custom kernels that correspond more closely to particular Green’s functions.
1.3 SVMs and Green’s Functions
Here, we briefly review the essential concepts of Support Vector Machines (SVMs) and quantum Green’s functions required for our subsequent discussions.
1.3.1 Key Elements of SVM Kernel Methods and Quantum Green Functions
Support Vector Machines [1, 2] aim to find a hyperplane that separates data into different classes (or fits a regression function) with maximum margin.
Common kernel functions include:
  • Radial Basis Function (RBF):
@@ 1.3.1 Key Elements of SVM Kernel Methods and Quantum Green Functions
  • Linear Kernel:
@@
  • Polynomial Kernel:
    $(\gamma\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+r)^{d}$ (3)
  • Sigmoid Kernel:
    $\tanh(\gamma\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+r)$ (4)
In quantum mechanics, the Green’s function $G(x,x^{\prime};E)$ is the inverse of the operator $\hat{H}-E$, where $\hat{H}$ is the Hamiltonian. Physically, $G$ describes how a disturbance at $x^{\prime}$ propagates to $x$. In imaginary time, the Euclidean Green’s function $G_{E}(x,x^{\prime};\tau)$ often appears in a Gaussian-like form, such as:
$G_{E}(x,x^{\prime};\tau)=\left(\frac{m}{2\pi\hbar\tau}\right)^{d/2}\exp\left(-\frac{m(x-x^{\prime})^{2}}{2\hbar\tau}\right)$ (5)
where $\tau$ is the imaginary time, $m$ is the mass of the particle, $d$ is the spatial dimension, $\hbar$ is the reduced Planck’s constant.
1.3.2 SVM Regression for Predicting Physical Quantities
Although Support Vector Machines are widely known as classification tools, they can also tackle real-valued predictions through Support Vector Regression (SVR). In this approach, the core optimization shifts from maximizing classification margin to minimizing regression error while preserving a margin-based structure in feature space.
Core Idea.
For each physical problem, we extract features $\mathbf{x}_{i}$ (e.g. wavevectors, angles, potential parameters) and associate them with continuous targets $y_{i}\in\mathbb{R}$ (such as conductivity or energy). SVR then fits a function
$f(\mathbf{x})\;=\;\sum_{j=1}^{N}(\alpha_{j}-\alpha_{j}^{*})\,K\bigl{(}\mathbf{x},\mathbf{x}_{j}\bigr{)}\;+\;b,$
where $K$ is a kernel function aligned with the system’s physical structure (e.g. polynomial for anharmonic expansions, RBF for Gaussian-like behavior, etc.).
We demonstrate how different kernels can address (1) conductivity with band-structure features, (2) scattering amplitudes via near-linear dependencies, (3) energy levels in an anharmonic oscillator, and (4) photonic or other periodic systems through custom kernels.
2 Mathematical and Formal Equivalence Between Kernel Functions and Green’s Functions
@@
In this section, we explore the mathematical equivalence between SVM kernel functions and quantum Green’s functions, highlighting their shared roles in mapping and operator inversion.
2.1 Comparing the Mathematical Forms
@@
The RBF kernel in SVMs, as defined in Eq. (1), and the Euclidean Green’s function, as defined in Eq. (5), both exhibit Gaussian dependence on the squared distance between points. By identifying the parameters appropriately, we observe that:
@@
where $\gamma$ is the kernel parameter in the RBF kernel, $m$ is the mass of the particle, $\hbar$ is the reduced Planck constant, and $\tau$ is the imaginary time.
@@
On the other hand, mapping $\phi(\mathbf{x})$ in SVMs can be seen as analogous to the quantum state $|\psi(\mathbf{x})\rangle$ in the Hilbert space. In both cases, we have vectors in high-dimensional spaces, in which inner products play a crucial role. In SVMs:
$K(\mathbf{x},\mathbf{x}^{\prime})=\langle\phi(\mathbf{x}),\phi(\mathbf{x}^{\prime})\rangle_{\mathcal{H}}.$ (8)
In quantum mechanics:
$G(\mathbf{x},\mathbf{x}^{\prime};t)=\langle\mathbf{x}e^{-i\hat{H}^{\prime}t/\hbar}|e^{i\hat{H}^{\prime\prime}t/\hbar}\mathbf{x}^{\prime}\rangle.$
$=\langle\psi(\mathbf{x,t})|\psi(\mathbf{x^{\prime},t})\rangle$ (9)
Here, $e^{-i\hat{H}^{\prime}t/\hbar}$ and $e^{i\hat{H}^{\prime\prime}t/\hbar}$ represent the evolution operators applied to the bra and ket states, respectively, under potentially different Hamiltonians $\hat{H}^{\prime}$ and $\hat{H}^{\prime\prime}$. This structure mirrors the kernel formulation in SVMs, where the transformation $\phi$ maps input data into a higher-dimensional feature space.
Furthermore, the feature mapping in SVMs, which maps data points into high-dimensional feature space, is analogous to the construction of the Green’s function in quantum mechanics, where the quantum states undergo time evolution. In both cases, these transformations provide a mathematical framework for evaluating the inner product, capturing the similarity or interaction between two states or data points in their respective spaces.
The inner product between states $|\psi\rangle$ and $|\phi\rangle$ in quantum mechanics is:
@@
This formulation highlights how quantum states interact in Hilbert space, with the integral serving as a measure of their overlap. Similarly, in SVMs, the kernel function represents the inner product in the high-dimensional feature space, capturing the similarity between data points after being mapped via the transformation $\phi(\mathbf{x})$.
In both SVMs and quantum mechanics, computations in high-dimensional spaces are handled implicitly, without explicit mapping taking place:
  • In SVMs, the feature mapping $\phi(\mathbf{x})$ into a high-dimensional space is not explicitly computed; instead, computations rely on the kernel function $K(\mathbf{x},\mathbf{x}^{\prime})$, which directly evaluates the inner product in the transformed space.
  • In quantum mechanics, the path integral formulation evaluates quantum amplitudes by summing over all possible paths that a system can take, thereby encapsulating the contributions from an infinite-dimensional space of paths without explicitly enumerating each path.
This implicit handling of high-dimensional spaces highlights a shared mathematical strategy in both fields, enabling efficient computations that may otherwise have been intractable.
2.2 Formal Equivalence Proof Through Operator Inversion, Mercer’s Theorem and Regularization
To claim that an SVM kernel exactly corresponds to a Green’s function, one must ensure that the operators governing both objects match in their spectral properties, boundary conditions, and self-adjointness or positivity, in addition to showing similarity in their functional form. To that end, we consider the issue from the viewpoint of operator inversion, where we can systematically incorporate conditions such as Mercer’s theorem (positive semi-definiteness), PDE boundary constraints, and eigenfunction expansions. Verifying all these aspects in tandem provides a stronger assertion of equivalence.
In quantum mechanics and PDE theory, the Green’s function is defined as the inverse (or resolvent) of the operator governing the system, e.g.,
$G=(\hat{H}-E)^{-1}\quad\text{or}\quad G=(\hat{L}-\lambda)^{-1},$
which describes how excitations or waves propagate under a given physical operator. Similarly, in SVMs the kernel matrix (or Gram matrix) plays a role analogous to an operator whose inversion (or regularized inversion) yields the support vector coefficients. This view of both the Green’s function and the kernel matrix as objects emerging from the inversion of an operator unifies the two frameworks, as discussed in more detail in Section 2.2.1.

### v2

@@ Abstract
We establish a mathematical equivalence between Support Vector Machine (SVM) kernel functions and quantum propagators represented by time-dependent Green’s functions, which has remained largely unexplored. We demonstrate that many common SVM kernels correspond naturally to Green’s functions via operator inversion theory. The sigmoid kernel does not always satisfy Mercer’s theorem, and therefore the corresponding Green’s function may also fail to perform optimally. We further introduce a Kernel Polynomial Method (KPM) for designing customized kernels that align with Green’s functions. Our numerical experiments confirm that employing positive-semidefinite kernels that correspond to Green’s functions significantly improves predictive accuracy of SVM models in physical systems.
1 Introduction
In recent years, machine learning methodologies have increasingly impacted physical sciences, providing powerful tools for data-driven modeling, pattern recognition, and even analytical insights into complex systems. A prominent example is the Support Vector Machine (SVM), a robust algorithm widely utilized for classification and regression tasks [1, 2]. Central to SVMs is the kernel trick, which implicitly maps input data into high-dimensional feature spaces, enabling the effective modeling of nonlinear relationships [3].
In parallel, quantum physics extensively employs Green’s functions to characterize particle propagation and system responses to external perturbations [4]. Quantum propagators, as time-dependent Green’s functions, describe how perturbations propagate through space-time. A notable illustration is the many-body Kubo formula for electrical conductivity, which fundamentally relies on quantum propagators to analyze transport properties.
Despite originating from distinct fields, SVM kernel functions and Green’s functions exhibit compelling mathematical parallels, yet this relationship has remained underexplored, particularly regarding their potential to enhance the prediction of physical quantities using machine learning. Specifically, the connection between common SVM kernels and Green’s functions arises naturally through the operator inversion, a key operation in quantum theory. Recognizing this relationship suggests a novel approach: by selecting kernel functions aligned with Green’s functions’ mathematical properties, we can significantly improve the predictive capabilities of SVM-based models in physics.
@@ 1 Introduction
Among the most recent approaches, we have Dean et al. [8] who use kernels to describe quantum correlations of $N$ non-interacting spinless fermions in their ground state. Then, they provide a method to compute the kernel in terms of the Green’s function for the corresponding single particle Schrödinger equation. On the other hand, Gin et al. [9] for example use deep learning to learn the Green’s function kernels with the purpose of solving non-linear differential equations. Li et al. [10] do the same for linear partial differential equations. However, the literature in the field is limited, and mostly constrained to applications in mathematics. To the best of our knowledge, all the previous studies were concerned with identifying equivalence relations between single kernels for a particular problem and Green’s functions.
1.1 Contributions
This paper systematically expands on these insights by providing a comprehensive study of the broader correspondence between kernel methods and Green’s functions within a machine learning framework applied to quantum physics. We speak of the equivalence between SVM kernels and quantum propagators as Green’s functions are related to quantum propagators via Laplace or Fourier transform, as discussed in Section 2.1.2.
Specifically, our contributions include:
  1. 1.
    Establishing a mathematical correspondence between commonly used SVM kernels and Green’s functions. We consider the mathematical form, the analogy from operator inversion, and spectral and eigenfunction decompositions.
  2. 2.
    Introduction of a Kernel Polynomial Method (KPM) for designing customized kernels that closely approximate specific Green’s functions.
  3. 3.
    Numerical experiments using kernels that align with Green’s functions to validate our theoretical findings. The experiments involve calculations of electrical conductivity, scattering amplitudes, energy spectra of anharmonic oscillators, and phonon dispersion in photonic crystals. All of the experiments demonstrate an enhanced predictive accuracy.
  4. 4.
    A strict equivalence between SVM kernels and Green’s functions would need to involve the respective boundary conditions. However, boundary conditions are typically hard to compute and so we don’t attempt this in the present work. Instead, based on the outcomes of our numerical experiments, we posit that considering them may not be necessary to obtain good predictive accuracy of kernels as long as they are aligned with a Green’s function.
The remainder of this paper is organized as follows: Section 2 introduces the concepts of SVMs and Green’s functions, and relates Green’s functions to quantum propagators. Section 3 reviews the theoretical foundations of SVM kernels and quantum propagators. Section 4 presents our numerical results. Section 5 details our methodological approach (including the implementation of KPM), while Section 6 concludes with a discussion of our findings and prospects for future research.
2 SVMs and Green’s Functions
In this section, we briefly review the essential concepts of SVMs and Green’s functions required for subsequent discussions.
2.1 SVM kernel methods and Green functions
SVMs [1, 2] are fundamental machine learning models that aim to construct a hyperplane that maximizes the margin between different classes in classification tasks or minimizes the error in regression tasks. However, when data is not linearly separable in its original feature space, kernel methods provide a powerful technique to implicitly map the data into a higher-dimensional space where a linear separation becomes feasible.
The core idea of kernel methods is to replace the standard inner product $\mathbf{x}^{\top}\mathbf{x}^{\prime}$ with a kernel function $K(\mathbf{x},\mathbf{x}^{\prime})$, which implicitly defines an inner product in a higher-dimensional (or even infinite-dimensional) feature space:
$K(\mathbf{x},\mathbf{x}^{\prime})=\phi(\mathbf{x})^{\top}\phi(\mathbf{x}^{\prime})$ (1)
where $\phi(\cdot)$ is a feature mapping function.
2.1.1 Common kernel functions
Several widely used kernel functions provide different properties suitable for various types of data structures and learning tasks. The four fundamental kernel functions are:
  • The radial basis function (RBF) kernel
@@ 2.1.1 Common kernel functions
    is particularly effective for capturing local structures in data.
  • The linear kernel
@@
    is suitable for linearly separable data.
  • The polynomial kernel
    $K(\mathbf{x},\mathbf{x}^{\prime})=(\gamma\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+r)^{d}$ (4)
    tends to be computationally efficient in modeling nonlinear relationships. It is used typically in cases where the linear kernel is too simple and the RBF kernel may overfit.
  • The sigmoid kernel
    $K(\mathbf{x},\mathbf{x}^{\prime})=\tanh(\gamma\,\mathbf{x}^{\top}\mathbf{x}^{\prime}+r)$ (5)
    may be used on mildly nonlinear datasets. It may fail to be positive semidefinite under certain parameter settings.
Besides these, some less known kernel functions include:
  • The Laplacian kernel
    $K(\mathbf{x},\mathbf{x}^{\prime})=\exp\!\bigl{(}-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|}{\sigma}\bigr{)},$ (6)
    a variant of the RBF kernel that is particularly effective on data that exhibits sharp decision boundaries.
  • The exponential kernel
    $K(\mathbf{x},\mathbf{x}^{\prime})=\exp\!\bigl{(}-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|}{2\sigma^{2}}\bigr{)}$ (7)
    which is similar to both the Laplacian and RBF kernels but offers smoother decision boundaries.
  • The Bessel kernel
    $K(\mathbf{x},\mathbf{x}^{\prime})=\frac{J_{\nu}(\|\mathbf{x}-\mathbf{x}^{\prime}\|)}{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{\nu}}$ (8)
    which is used in specialized applications, such as wave propagation modeling.
  • The ANOVA kernel
    $K(\mathbf{x},\mathbf{x}^{\prime})=\sum_{i=1}^{n}\exp(-\gamma(x_{i}-x^{\prime}_{i})^{2})$ (9)
    used in modeling interactions in datasets where variables exhibit structured dependencies.
In this paper, we focus on the four fundamental kernel functions: RBF, linear, polynomial, and sigmoid. These kernels are the primary objects of study in Section 4, where we analyze their properties and relationships to Green’s functions. The other kernel functions listed above are considered either as natural extensions or as part of Section 5, where we explore custom kernels and their generation using the KPM approach.
2.1.2 Quantum propagators and Green’s functions
In quantum mechanics and field theory, quantum propagators and Green’s functions are connected concepts used to describe the evolution of quantum states and the response of systems to external influences.
The quantum propagator, or Feynman propagator, describes how a quantum state evolves from an initial state $x$ to a final state $x^{\prime}$ over time $t$. It is given by the transition amplitude:
$K(x^{\prime},t;x,0)=\langle x^{\prime}|e^{-iHt/\hbar}|x\rangle$ (10)
where $K(x^{\prime},t;x,0)$ is the propagator, $H$ is the Hamiltonian of the system, and $e^{-iHt/\hbar}$ is the time evolution operator.
The propagator satisfies the Schrödinger equation and describes the probability amplitude of a particle moving from $x$ to $x^{\prime}$.
The path integral formulation expresses the propagator as:
$K(x^{\prime},t;x,0)=\int\mathcal{D}[x]e^{iS[x]/\hbar}$ (11)
where $S[x]$ is the action in the Lagrangian formulation.
A Green’s function is a mathematical tool used to solve differential equations, especially those involving operators such as the Hamiltonian $H$:
$(H-E)G(x,x^{\prime})=\delta(x-x^{\prime})$ (12)
where $G(x,x^{\prime})$ is the Green’s function, $H$ is the Hamiltonian operator, $E$ is the energy, and $\delta(x-x^{\prime})$ is the Dirac delta function.
The connection between the quantum propagator and the Green’s function is given by the following equality:
[…]
which shows that the Green’s function is the Laplace or Fourier transform of the propagator.
Alternatively, in momentum space:
$G(k,E)=\frac{1}{E-H+i\epsilon}$ (14)
where $G(k,E)$ is the momentum-space Green’s function, and $i\epsilon$ ensures causality and defines time ordering.
From the above it follows that the propagator describes the time evolution of a quantum system, while the Green’s function solves for quantum states and characterizes the response of the system. As quantum propagators and Green’s functions are related through the Fourier/Laplace transform, and can therefore be seen as dual perspectives on the same matter, subsequent discussions will be framed in terms of Green’s functions.
In relativistic quantum field theory (QFT), the Feynman propagator is a Green’s function for the Klein-Gordon or Dirac equation:
$(\partial^{\mu}\partial_{\mu}+m^{2})G_{F}(x,x^{\prime})=-\delta^{4}(x-x^{\prime})$ (15)
which governs the behavior of quantum fields.
3 Mathematical equivalence between SVM kernels and Green’s functions
@@
In this section, we explore the mathematical equivalence between SVM kernel functions and Green’s functions, highlighting their shared roles in mapping and operator inversion.
3.1 Comparison of mathematical forms
In imaginary time, the Euclidean Green’s function $G_{E}(\mathbf{x},\mathbf{x^{\prime}};\tau)$ often takes a Gaussian-like form:
$G_{E}(\mathbf{x},\mathbf{x^{\prime}};\tau)=\left(\frac{m}{2\pi\hbar\tau}\right)^{d/2}\exp\left(-\frac{m(\mathbf{x}-\mathbf{x^{\prime}})^{2}}{2\hbar\tau}\right)$ (16)
where $\tau$ is the imaginary time, $m$ is the mass of the particle, $d$ is the spatial dimension, and $\hbar$ is the reduced Planck’s constant.
@@ 3.1 Comparison of mathematical forms
The RBF kernel in SVMs is defined as in Eq. (2). By comparing Eqs. (16) and (2) we find that both exhibit a Gaussian-like structure, with the Euclidean Green’s function describing diffusion in imaginary time and the RBF kernel defining a similarity measure in feature space. This analogy suggests that certain kernel-based methods in machine learning share mathematical foundations with Green’s functions in physics, and both exhibit Gaussian dependence on the squared distance between points. By identifying the parameters appropriately, we observe that:
@@
where $\gamma$ is the kernel parameter in the RBF kernel given in Eq. (2).
@@
On the other hand, mapping $\phi(\mathbf{x})$ in SVMs can be seen as analogous to the quantum state $|\psi(\mathbf{x})\rangle$ in the Hilbert space. In both cases these are vectors in high-dimensional spaces, in which inner products play a crucial role. In SVMs, we have
$K(\mathbf{x},\mathbf{x}^{\prime})=\langle\phi(\mathbf{x}),\phi(\mathbf{x}^{\prime})\rangle_{\mathcal{H}},$ (19)
while in quantum mechanics, we have
$G(\mathbf{x},\mathbf{x}^{\prime};t)=\langle\psi(\mathbf{x})e^{-i\hat{H}^{\prime}t/\hbar}|e^{i\hat{H}^{\prime\prime}t/\hbar}\psi(\mathbf{x^{\prime}})\rangle.$
$=\langle\psi(\mathbf{x,t})|\psi(\mathbf{x^{\prime},t})\rangle.$ (20)
Here, $e^{-i\hat{H}^{\prime}t/\hbar}$ and $e^{i\hat{H}^{\prime\prime}t/\hbar}$ represent the evolution operators applied to the bra and ket states, respectively, under potentially different Hamiltonians $\hat{H}^{\prime}$ and $\hat{H}^{\prime\prime}$. This structure mirrors the kernel formulation in SVMs, where the transformation $\phi$ maps input data into a higher-dimensional feature space. This mapping is analogous to the construction of the Green’s function in quantum mechanics, where the quantum states undergo time evolution. In both cases, these transformations provide a mathematical framework for evaluating the inner product, thereby capturing the similarity or interaction between two states or data points in their respective spaces. The inner product between states $|\psi\rangle$ and $|\phi\rangle$ in quantum mechanics is given by
@@
and it serves as a measure of their overlap. Similarly, the kernel function in SVMs captures the similarity between data points after being mapped by $\phi(\mathbf{x})$.
Computations of both SVMs and quantum mechanics are handled implicitly in high-dimensional spaces, without explicit mapping taking place, but there is a major difference. In SVMs, the feature mapping $\phi(\mathbf{x})$ into a high-dimensional space is not explicitly computed; instead, computations rely on the kernel function $K(\mathbf{x},\mathbf{x}^{\prime})$, which directly evaluates the inner product in the transformed space.
In quantum mechanics, we must explicitly describe the time evolution of the wave function, rather than merely obtaining the final inner product result.
3.2 Other forms of equivalence
To claim that an SVM kernel exactly corresponds to a Green’s function, one must ensure that the operators governing both objects match in their spectral properties, boundary conditions, and self-adjointness or positivity, in addition to showing similarity in their functional form.
In quantum mechanics and the theory of partial differential equations (PDE), Green’s function is defined as the inverse (or resolvent) of the operator governing the system, e.g.,
$G=(\hat{H}-E)^{-1}\quad\text{or}\quad G=(\hat{L}-\lambda)^{-1},$ (22)
where:
  • $\hat{H}$ is the Hamiltonian operator, describing the total energy of a quantum system.
  • $E$ is the energy eigenvalue, satisfying the eigenvalue equation $\hat{H}|\psi\rangle=E|\psi\rangle$.
  • $\hat{L}$ is a differential operator in PDE theory, such as the Laplacian $\nabla^{2}$, governing wave propagation or diffusion.
  • $\lambda$ is an eigenvalue of $\hat{L}$, associated with solutions of the PDE.
A Green’s function describes how excitations or waves propagate under a given physical operator. Similarly, in SVMs, the kernel matrix (or Gram matrix) plays a role analogous to an operator whose inversion yields the support vector coefficients. This view of both Green’s function and the kernel matrix as objects emerging from the inversion of an operator unifies the two frameworks, as discussed in more detail in Section 3.2.1.